OpenStudy (anonymous):

A cereal box is 8 inches long, 3 inches wide, and 12 inches tall. (a) Volume? (b) Surface area?

OpenStudy (anonymous):

Volume = Length x Width x Height = 8 x 3 x 12 if L = Length, W = Width, and H = Height, Surface Area = 2LW + 2LH + 2WH = 2(8)(3) + 2(8)(12) + 2(3)(12)

OpenStudy (anonymous):

Volume = 8*3*12 = 288

OpenStudy (anonymous):

Volume = 8*3*12 = 288